OpenStudy (anonymous):

Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(1/3) is undefined.

jhonyy9 (jhonyy9):

so where is f(x)=0 where is this function f(x) ?

OpenStudy (anonymous):

oh sorry. f(x) = x - 1

jhonyy9 (jhonyy9):

so than f(1/3)= (1/3) -1 =(1-3)/3 = -2/3 so this not is undefined just if there in your mathbooke from this function f(x)=x-1 being again anything property added - do you see again there anything from f(x) ?

OpenStudy (anonymous):

well see i thought the same thing. But my paper is asking me why is that one undefined and i dont understand what you mean by your question.

jhonyy9 (jhonyy9):

so for examnple if there being ,has wrote that f(x) is always positiv so than f(1/3) will be undefined ok ?

OpenStudy (anonymous):

Well in the beginning it asked me to simplify (3x^2 - 4x + 1)/(3x - 1).. Then it asks me what you already know.

jhonyy9 (jhonyy9):

or if x need to be integer so than f(x) will be undefined - i think i think that f(1/3) can being considered undefined if like property of f(x) ok so now i understand why f(1/3) is undefined

jhonyy9 (jhonyy9):

so check it newly what you have there like function ? so you see that there is a fraction

jhonyy9 (jhonyy9):

and how do you know it when a fraction is undefined ?

jhonyy9 (jhonyy9):

so for denominator there is 3x-1 ,and if you replace,substitute 1/3 in place of x what will get ?

jhonyy9 (jhonyy9):

hope so much that now you understand it right sure easy yes ?

OpenStudy (anonymous):

I'm not sure when youknow whether or not a funciton can be undefined or not. When you substitute it into the equation f(x) = x - 1, you have 0.

jhonyy9 (jhonyy9):

so because there is a fraction with denominator 3x-1 yes ?

jhonyy9 (jhonyy9):

so and because 3*(1/3) -1 =0 so for this case because a fraction not can being with denominator zero so for this case will be f(1/3) undefined ok ?

jhonyy9 (jhonyy9):

in the begining step check your exersice ok ?

OpenStudy (anonymous):

ok i think i am starting to get it. tahnk you so much!!

jhonyy9 (jhonyy9):

yw good luck bye